OpenStudy (tiffclifford13):

I'm in desperate need of help Graph the function using the techniques of shifting, compressing, stretching, and/or reflecting. Start with the graph of the basic function and show all stages. f(x) = 3(x - 2)^2 + 1

OpenStudy (danjs):

The original function is the parabola with vertex on (0,0) f(x) = x^2

OpenStudy (danjs):

For f(x) = a*f(x - b) + c the a,b,and c values change the function. b - shift in horizontal direction c - shift in vertical direction a - squeeze or stretch the graph vertically

OpenStudy (danjs):

\[f(x) = 3*(x-2)^2 + 1\] horizontal shift by 2, moves the +x direction 2 vertical shift by 1, moves the +y direction 1 stretches the graph vertically by a factor of 3

OpenStudy (tiffclifford13):

Let me start by saying thank you so much for that I appreciate it but I'm still confused of how I would write that like when you're talking about the +x direction 2 and +y direction 1 could you show me please just like a rough example of it?

OpenStudy (mathmate):

hint: read all about it here: https://www.mathsisfun.com/sets/function-transformations.html

OpenStudy (mathmale):

To address your specific question: Suppose you start with the graph of y=x^2. Were you to translate (move or shift) this graph "a" units to the right, the function would become y=(x-a)^2. "a" units to the left, the function would become y=(x+a)^2.